Problem 2. (4 pt) Use the Laplace transform to solve the following initial-value problem: y + 4y = 13e3t, y(0) = 2, y(0) =

Please, I need help.
Answer #1

y+4y= 138, y(0) = 2, yO) =-3 using Laplace transform L{ y + 4yY= 13 lety ŠY(5) – syco) – y(0) +49(5) - Š¥(5)-25+3+44(S) = 13*13263+4) = 18.15-3) (3+4)- Gaze :16) = 53 23 ste minta to #4 Yis) = 1 t s S-3 Š +2 3+2 verse taking inverse Laplace transf